OpenStudy (anonymous):

what is the least common multiple of 6 and 21?

OpenStudy (anonymous):

I don't remember how to do this?

OpenStudy (anonymous):

multiply them until they equal the same number and divide it to the simpliest number

OpenStudy (dangerousjesse):

Find the prime factorization of your integers first

OpenStudy (dangerousjesse):

\(\large 6=2\times 3\)

OpenStudy (dangerousjesse):

\(\large 21=7\times 3\)

OpenStudy (dangerousjesse):

Then find the largest power of each prime factor

OpenStudy (dangerousjesse):

So the largest power of 2 that appears in factorization is \(\large 2^1\)

OpenStudy (dangerousjesse):

For three it's \(\large 3^1\)

OpenStudy (dangerousjesse):

And for seven it's \(\large 7^1\)

OpenStudy (dangerousjesse):

So your LCM= \(\large 2^1\times 3^1\times 7^1\)

OpenStudy (dangerousjesse):

LCM(6,21)=42 :)

OpenStudy (dangerousjesse):

Now do you understand?

OpenStudy (anonymous):

I don't know why I couldn't reply before??? that was weird. but no I really don't get this as ive been out of school 10 years